Operations Report – February 24th

SOL: 0

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need
replacement on generator. A contractor has been arranged to make these
repairs within a few days.

Generator (hours run): Turned off at 1430 (24th Feb), Turned on at
1903 (24th Feb)

Solar— SOC % – Turned off Gen (24th Feb) 75%, Turned on Gen (24th Feb) 100%

Diesel Reading – ~ 75%

Propane Reading – ~ 26%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – around 76%, ~ 419 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 152,9 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01415558 units

Toilet tank emptied: Yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Used

Hours: 51.3 hours

Beginning charge: 100%

Ending charge: 74%

Currently charging: Yes

Curiosity rover used: Used

Hours: 75.5 hours

Beginning charge: 100%

Ending charge: 75.5%

Currently charging: Yes

Spirit rover used: Used

Hours: 78.3 hours

Beginning charge: 100%

Ending charge: 66%

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Used to bring back crew 205 to Grand Junction

Crew Car used and why, where? Used to bring back crew 205 to Grand Junction

General notes and comments: In the Hab’s upper deck, the bulb from one
of the lamps (on the sink side) is menacing to fall, one screw is not
holding anymore. Also this lamp is holding onto the electrical cable
rather than its theoretical link to the ceiling.

Summary of internet: Nothing to report

Summary of suits and radios: New inventory of the radios this morning, five of them have minor to major issues. Going to test them again tomorrow after full charging.

Summary of Hab operations: Nothing to report

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Propane tank
needs to be filled rapidly, how do we proceed ?